Ask your own question, for FREE!
Statistics 17 Online
OpenStudy (anonymous):

A die is rolled three times. Find the chance that the face with 4 spots appears at least one.

OpenStudy (anonymous):

Let A=probability getting 4=1/6 B=probability not getting4=1-1/6=5/6 In all three rolls of die P(not getting 4)=B*B*B=5/6*5/6*5/6=125/216 p atleast one 4=1-125/216=(216-125)/216=91/216

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
1 day ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
1 day ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!